HETAV PATEL

Grade 11,

two masses M1=1 kg and M2=2 kg are connected by a mass less spring. a force of 4N acts on the 2kg mass. what will be accn of M2 if accn of M1 is 1m/s^2

two masses M1=1 kg and M2=2 kg are connected by a mass less spring. a force of 4N acts on the 2kg mass. what will be accn of M2 if accn of M1 is 1m/s^2

Grade:11

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Zahid,

 FOr M1 : eq of motion -       M1 a1 = kx or kx = 1 N

For M2 : eq of motion  -     F - kx = M2a2     

                                        or  4 - 1 = 2 a2

                                              a2 = 3/2 m/s2

 

 Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil pannikar

IIT Bombay

 

 

suryakanth AskiitiansExpert-IITB
13 years ago

Dear Zahid,

considering the free body diagram of M2,forces in it are,

F-Fspring = M2*a

on mass M1 ,forces in horizintal direction are

Fspring = M1*a1

Given s1 = 1 m/s^2

Hence Fspring = 1 N

M2*s2 = Fgiven - F spring

M2*s2 = 4 - 1 = 3 N

Hence accelaration of the MASS M2 = 1.5 m/s^2

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB